OpenStudy (anonymous):

f(x)=3x^3-2x^2+5x-5 find f'(2)

OpenStudy (anonymous):

33

OpenStudy (anonymous):

f'(x)=3*3x^2-2*2x+5-0 f'(2)=3*3*2^2-2*2*2+5 =9*4-8+5=36-8+5=33

OpenStudy (anonymous):

in order to solve this, plug in (2) for every X